# Large-order perturbation theory of linear eigenvalue problems

### v1

@@ 5 Example 4: Divergence driven by two singularities
Our final example is chosen to illustrate that the divergence of the eigenvalue can be driven by more than one singularity in the outer solution, leading to more exotic behaviour. This is exactly what happens in the model in [7] concerning quasinormal modes of Reissner-Normström de Sitter black holes. The form of this divergence is more difficult to pick up with other methods, and the interaction between two singularities makes it difficult to guess the form of the divergence from numerical calculations of the leading terms in the series.
@@ 5.2 Outer region
The next step is to determin $\Lambda_{1}$ and $\Lambda_{2}$ through matching with inner regions near $X=-c\pm\mathrm{i}b$.
@@ 6 Conclusion
However, the late term approximation of this outer expansion also non-uniform, now not as $\epsilon\rightarrow 0$ but as $n\rightarrow\infty$. By introducing a local variable in the equation for the late terms of the outer expansion, a new inner expansion is generated in which the two parts of the diverence become coupled, and the eigenvalue is determined.
@@ Acknowledgements
I would like to acknowledge the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme ‘Applicable resurgent asymptotics: towards a universal theory’, where part of the work on Examples 1 and 3 was undertaken, as well as useful discussions with many of the participants. I would like to single out in particular Ines Incieto, who among other things introduced me to reference [7] during this programme. This programme was supported by EPSRC grant no. EP/R014604/1.

### v2

@@ 5 Example 4: Divergence driven by two singularities
Our final example is chosen to illustrate that the divergence of the eigenvalue can be driven by more than one singularity in the outer solution, leading to more exotic behaviour. This is exactly what happens in the model in [7] concerning quasinormal modes of Reissner-Nordström-de Sitter black holes. The form of this divergence is more difficult to pick up with other methods, and the interaction between two singularities makes it difficult to guess the form of the divergence from numerical calculations of the leading terms in the series.
@@ 5.2 Outer region
The next step is to determine $\Lambda_{1}$ and $\Lambda_{2}$ through matching with inner regions near $X=-c\pm\mathrm{i}b$.
@@ 6 Conclusion
However, the late term approximation of this outer expansion also non-uniform, now not as $\epsilon\rightarrow 0$ but as $n\rightarrow\infty$. By introducing a local variable in the equation for the late terms of the outer expansion, a new inner expansion is generated in which the two parts of the divergence become coupled, and the eigenvalue is determined.
@@ Acknowledgements
I would like to acknowledge the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme ‘Applicable resurgent asymptotics: towards a universal theory’, where part of the work on Examples 1 and 3 was undertaken, as well as useful discussions with many of the participants. I would like to single out in particular Ines Aniceto, who among other things introduced me to reference [7] during this programme. This programme was supported by EPSRC grant no. EP/R014604/1.